• A 150Ω resistor is connected in series to a1kΩ resistor and to a battery of 127V. Find the effective resistance and current in the circuit.
1
Expert's answer
2021-11-09T16:36:13-0500
"R=1000+150=1150\Omega\\
I=\frac{127}{1150}=0.11\ A"